OpenStudy (anonymous):

I am still practicing integration by parts. If I have mistakes please catch me on it...

OpenStudy (anonymous):

the problem is:\[\int\limits_{ }^{ }\sin^{-1}x~dx\]

OpenStudy (anonymous):

I am going to make the second function g(x)=1, and f(x)=sin^(-1)(x). (this is how I learned this) So I will need to find the derivative of sin^(-1)(x).

OpenStudy (anonymous):

\[y=\sin^{-1}(x)\]\[\sin(y)=x\]differentiating.\[y'~\cos(y)=1\]\[y'=\sec(y)\]SO,\[y'=\frac{1}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

no I can proceed with integration. \[\int\limits_{ }^{ } \sin^{-1}(x)~dx=x \sin^{-1}(x)- \int\limits_{ }^{ } \frac{x}{\sqrt{1-x^2}}~dx\]

OpenStudy (anonymous):

then I am going to do a u substitution. \[u=1-x^2\]

OpenStudy (anonymous):

\[-\frac{1}{2}du=x~dx\]

OpenStudy (anonymous):

\[\int\limits_{ }^{ } \sin^{-1}(x)~dx=x \sin^{-1}(x)+ \frac{1}{2}\int\limits_{ }^{ } \frac{1}{\sqrt{u}}~du\]

OpenStudy (anonymous):

then applying power rule: u^(-1/2) ---> 2sqrt{u} and then we sub back the x.

OpenStudy (anonymous):

\[\int\limits_{ }^{ } \sin^{-1}(x)~dx=x \sin^{-1}(x)+\sqrt{1-x^2}+C\]

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

this is calculus. I am practicing integration by parts.

OpenStudy (anonymous):

can you check it please?

OpenStudy (anonymous):

you didn't learn this yet?

OpenStudy (anonymous):

oh, I see. I'll wait for someone to come by and verify if I am right or wrong.

OpenStudy (anonymous):

@eliassaab can you check my problem?

OpenStudy (watchmath):

it's correct

OpenStudy (anonymous):

ty!